Define the unemployment rate as UR = U/(M + U) where U is the number of people unemployed in the manufacturing sector, and M is the number of people employed in the manufacturing sector. Then the expected wage in manufacturing can be written as E(WM) = 400(1 – UR).
C Suppose the economy starts out with all 100 workers in agriculture, and then 50 people move from agriculture to the manufacturing sector. Solve for the expected wage in manufacturing, assuming there are 50 people in the manufacturing sector and a legislated minimum wage of $400.
D What is the wage in agriculture with 50 people in the agricultural sector?
E Will more people in agriculture want to move to the manufacturing sector? Why
1
Expert's answer
2019-04-18T09:33:13-0400

C. UR = 0/50 = 0,

E(WM) = 400*(1 - 0) = $400.

D. The wage in agriculture with 50 people in the agricultural sector is not less than $400 too because there is a legislated minimum wage of $400.

E. People in agriculture will not want to move to the manufacturing sector, because the wage is the same.